44. Prove that if a and b are different integers, then there exist infinitely 

many positive integers n such that a+n and b+n are relatively prime. 


1
Expert's answer
2022-06-16T15:00:46-0400

Let "a" and "b" be two different integers. Assume for instance "a<b", and let "n=(b-a)k+1-a". For "k" sufficiently large "n" will be positive integer. We have "a+n=(b-a)k+1", "b+n=(b-a)(k+1)+1", hence "a+n" and "b+n" will be positive integers. If we had "d|a+n" and "d|b+n", we would have "d|a-b", and, in view "d|a+n", also "d|1", which implies that "d=1", Thus "(a+n,b+n)=1".